# Leveraging Deep Generative model for computational protein design and optimization

### v1

@@ 2 Computational Protein Design
Machine learning and deep learning based design:Machine learning based backbone generation models developed recently trained with structures from the PDB are used to generate de novo protein structures. For example, a generative adversarial network(GAN) based method [anand2019fully] can create protein structures by generating pairwise distance maps and decode with a pre-trained downstream neural networks. Another autoencoder base model[eguchi2022ig] focused on immunoglobulins can generate immunoproteins by decoding latent representations. More recently, a model developed can build distance map by iterative refinement and network hallucination[anishchenko2021novo] and repurposed a structure prediction model[yang2020improved] to generate three dimensional protein structures.A recent study that utilized this model demonstrated its ability toe facilitate the design novel enzymes with competitive enzymatic activity[yeh2023novo]. On the other hand, a slew of diffusion based models [watson2023novo, lin2023generating, wu2024protein, ingraham2023illuminating] emerged for structure biology tasks aimed to address challenges presented by previous methods, specifically the ability to flexibly generate diverse and high-quality protein structures. In addition to successfully designing novel monomeric proteins, protein binders and, enzyme active site scaffolds[watson2023novo] with experimental characterization. Despite its success, there are still drawbacks from diffusion based structure models which often demand large amount of compute and hard to incorporate physical constraints into the generative process. It is also worth noting that the success of functional protein design still heavily relies on extensive laboratory screening.
@@
Deep Learning Methods:More recent approaches use deep learning for sequence design, demonstrating superior performance in both computational tests and wet experiments[ingraham2019generative, dauparas2022robust, li2014direct, anand2022protein]. These methods can capture complex, non-linear relationships in protein structure and sequence, and have the ability to learn from large datasets of known protein structures. They are potentially more robust to variations in target structures. However, deep learning methods are often difficult to interpret the basis of their predictions. Additionally, they may struggle with novel protein folds not well-represented in training data[li2014direct, mcpartlon2022deep, gao2022pifold]. Besides, many application specific models are developed for sequence design aimed to enhanve performance in domain specific tasks such as antibody design[hoie2024antifold, dreyer2023inverse]. In this study, we will utilize our structure generative model and the iterative design framework in combination with a wide range of inverse folding models to improve the efficiency and robustness of computational protein design pipelines.
Structure-based de novo functional protein design, which involves creating proteins distinct from those exist in the nature based on biophysical principles, has achieved notable successes including the design of novel protein folds[huang2016novo] and the creation of de novo enzymes[rothlisberger2008kemp, yeh2023novo]. Protein redesign and optimization, which involves modifying existing proteins for enhanced stability, altered specificity, or new functions, has also seen substantial advancements [wijma2014computationally]. Although significant effort and progress has been made towards computationally designing novel structure and sequence to achieve new functional proteins. However, current successes in protein design often require extensive manual input from both biochemistry and computational experts, limiting the accessibility of advanced protein design tools to the broader scientific community and hindering their ability to achieve diverse research goals. This thesis will explore and discuss novel methods for developing a more accessible computational framework, aiming to allow easier access of advanced protein design tools and enable their efficient use for de novo protein design across the life science and computational communities.
@@ 3 Machine learning
KL divergence is non-negative according to the Jensen’s inequality. The log likelihood(evidence) is then lower-bounded by $E_{q_{\phi}}[\log(p_{\theta}(x,z))]-E_{q_{\phi}}[\log(q_{\phi}(z|x))]$, which is referred to as the Evidence Lower Bound(ELBO). We can estimate the gradient of the ELBO with a naive Monte Carlo estimator, however, it is often challenging due to its high variance[paisley2012variational]. Another important observation is that the evidence is a constant and the KL divergence is non-negative, maximizing the ELBO can be seen as simultaneously minimizing the KL divergence between the variational distribution and the true posterior and maximizing the marginalized data likelihood. We can then rewrite the ELBO as
@@ 3 Introduction
In recent years, progress in machine learning and deep learning research has contriburted to significant advances for protein modeling such as mutation effect estimation [frazer2021disease, meier2021language, riesselman2018deep], protein function prediction [lai2022accurate, gligorijevic2021structure], and structure prediction [jumper2021highly, baek2021accurate, wang2017accurate].These advancements have also played a part in aspects of the computational protein design problem. For fixed backbone sequence design, a series of deep learning methods have emerged to improve conventional energy based approaches [alford2017rosetta, khatib2011algorithm] by directly incorporating structural information using SE(3)-equivaraint frameworks [mcpartlon2022deep, jing2020learning, hsu2022learning]. An array of recent works have studied the use of generative models for structure generation[lin2021deep, anand2019fully, anand2018generative], however, these methods often generate topological constraints and rely on downstream tools for 3-dimensional structure determination. One of the major difficulties for direct coordinate generative modelling is properly accounting for the rotation and translation equivariance in the target conformation.
@@ 2 Ablation study on input features
In addition, we also compared the short, medium, long range contact accuracy where the predicted contacts are derived from the inter-residue distance of the generated 3-dimensional coordinates as shown in table 3.
@@ 3 Conformational decoy sampling for robust protein sequence design
We provide examples in Figure 4(A) where our designed sequence folded more successfully than the native sequence when MAS and template information is not available. Particularly, as shown for PDB:3G67, the native sequence is folded with RMSD = 36.26 while our best decoy designed sequence folded with RMSD = 2.56 with high confidence. We observed that our model can consistently improve designed quality regardless of the size of the target backbone. In Figure 4 (B), we deployed the design pipeline to a set of 211 backbone targets and plotted the TM-score of the predicted structure compared to native sequence(bottom), and sequence designed from the native backbone structure(top). Sequences designed from our pipeline outperform sequences designed from fixed native backbones in 856 out of the 1,016 tested targets(84%).
@@
To further demonstrate our deep structure generative model presents an edge in providing diverse backbone ensembles that can overcome limitations in inverse folding methods relying on single-shot native backbones. Fig. 6(A) illustrates an example where proteinMPNN temperature-based sequence sampling with a native backbone failed to generate structure-conforming sequences, while our structure ensemble approach yielded many successful designs. The advantages of our method is further evidenced in Fig. 6(B), which shows the distribution of sc-TM scores versus pLDDT across all tested targets. Notably, our conformer-based designs (red points) achieved a higher concentration of successful designs in the upper-right quadrant compared to single-shot native backbone templates (blue points). The bar chart below quantifies this advantage, showing that conformer-based designs consistently outperform native backbone designs across various sc-TM cutoffs, with approximately twice the success rate in many cases. Fig. 6(C) further supports this conclusion by showing that our conformer-based designs (blue points) consistently achieve higher pLDDT scores compared to native backbone designs across the majority of tested folds.
@@ 7 Discussion
In the following chapters, I will employ the structural generative model developed and explore de novo protein design and protein optimization through a iterative framwork and applying the model to challenging real-world protein design problems and validating designs experimentally. Extending the model to handle multi-chain protein complexes and further antibody design. However, other potential directions for future work includes integrating the structural generative model more tightly with sequence design algorithms, potentially in an end-to-end differentiable framework. Exploring the use of the learned structural embeddings for other tasks such as function prediction or protein-protein interaction modeling.
@@ 5 Iterative optimization algorithms
A set of randomly sampled sequence pool of length $L$ is initiated, for each iteration, we predetermine the number of templates we select for each design iteration and the number of decoys to generate for each template. Then we set the number of iteration to be $T$. In the design pipeline, Alphafold2 with single sequence mode was used as StructurePrediction and ProteinMPNN was used as the FixedBBDesign algorithm with 1 sequence generated for each decoy. RankPTM is a function that rank a set of prediction structure by their pTM scores.
@@
1:procedure MotifGroundedScaffold($Tempate_{init},T,N_{decoy},N_{template},MotifSites$)
@@
We initiate our design pipeline with a structure motif $Template_{init}$ of interest as well as the relative infilling among the segments of the motif, for each iteration, we predetermine the number of templates we select for each design iteration and the number of decoys to generate for each template. Then we set the number of iteration to be $T$. In the design pipeline, ESMFold is used as StructurePrediction and ProteinMPNN is used as the FixedBBDesign algorithm with 1 sequence generated for each decoy and the motif sequence were fixed in this step. RankMotifRMSD is a function that rank a set of prediction structure by their the RMSD w.r.t the input motif.
@@ 6 Data
We used the DIPS(Database of Interacting Protein Structures) dataset presented by [townshend2019end] to fine tune our structure generative model for binding sruface generation. All protein dimer complexes are process and we capped the size of the complex structure to 500 amino acid to account for memory usage. Then we computed the binding surface with a 10A radius inter-chain contact, that is, any residue that is within 10A to any residue of the other chain. During training, the binding interface of one component of the protein complex is randomly masked.
@@ 2 De novo design of DFHBI activated fluorescent protein
Figure 8: Experimental validation of de novo designed DFHBI-activated fluorescent $\beta$-barrels A) Fluorescence microscope image of de novo designed protein(Left) and the design template mFAP0. B) Fluorescent emission spectras from the lysates of 20 tested designs and the templates with reference buffer. All of the 20 designs were found with detectable fluorescent emission which represents a 100% design success rate. C) Fluorescent emission spectras of purified designed proteins and their respective original templates of 6czi(up) and 6czh(down). Although we did not observe increased fluorescent peak, we saw slight shit in the emission frequency compared to the design template. D)SDS–PAGE of purified designed protein with its respective design templates. All the designed proteins are observed to be smaller than the reference template. E) Protein thermal stability assay results(top), and protein yield comparison results(bottom). We see improvement of thermal stability on 6czh-based designs and comparable thermal performance in 6czi-based designs. For protein yield, all of the 6czh-based designs exhibited improvement over the design reference and 6czi-V has a significant increase over the reference.
We selected the top 20 designs from our computational pipeline for initial experimental screening. These designs were expressed in E. coli(See Methods for details) using standard protocols. Crude cell lysates were then subjected to a fluorescence assay in the presence and absence of DFHBI. This initial screen allowed us to rapidly assess the fluorescence activation of our designs and compare their performance to the design templates. Results from this initial screen revealed that 20 out of 20 designs showed detectable DFHBI-activated fluorescence a $100\%$ success rate8.B. All the tested designs have sequence similarity to the design template greater than $74\%$ with the lowest candidate to have $65\%$ sequence identity to the design template as shown in 8.A.
@@
Protein scaffolding is a crucial tool in protein engineering that involves transplanting functional motifs from one protein context to another, often more stable and versatile, protein scaffold[wang2022scaffolding]. This technique has numerous applications in biotechnology and therapeutics, allowing for the creation of proteins with desired functions in optimized structural contexts. Scaffolding can improve protein stability, enhance expression levels, and even modulate the activity of the transplanted motif. To trail our proposed iterative design framework’s ability to embed known functional motif into new protein scaffolds by iteratively co-optimizing structure and sequence around the dedicated motif of interest as seen in 9. We performed an excise on computationally scaffold the PD1/PD-L1 complex and validated our design with in silico binding and folding experiments.
@@
Programmed Cell Death Protein 1 (PD1) is a crucial immune checkpoint receptor that plays a significant role in regulating T-cell responses[sharpe2018diverse]. The interaction between PD1 and its ligands (PD-L1 and PD-L2) is a key target for cancer immunotherapy[sun2018regulation]. However, the use of native PD1 protein or its extracellular domain in therapeutic applications can be challenging due to stability and manufacturing considerations[ganesan2019comprehensive].In this study, we aimed to use our iterative design framework to scaffold the key binding motif of PD1 onto a more stable protein structure. Our goal was to create a novel protein that maintains PD1’s binding specificity and affinity for its ligands while potentially offering improved biophysical properties.
@@ 4 Structure based Antibody design via conditional CDR inpainting
In FigLABEL:fig:C2_R8 we observed that the two stage fine tuning significantly improved our model’s ability to accurately in paint the CDRs of interest. Our final fine tuned model $CoordVAE_{ab}$, demonstrates significant improvements across all three CDR regions compared to other models. It consistently achieves the highest TM-scores and pLDDT scores, indicating better overall structural similarity and prediction confidence. Notably, it also maintains the lowest RMSD values across all CDR types, suggesting more accurate reconstruction of the original structures. The performance gap is most significant for CDR3, which is often the most consequential CDR and the hardest to design. These results indicate that our model is exceptionally effective at generating accurate and structurally similar CDR regions, which is crucial for antibody design and engineering.

### v2

@@ 2 Computational Protein Design
Machine learning and deep learning based design:Machine learning based backbone generation models developed recently trained with structures from the PDB are used to generate de novo protein structures. For example, a generative adversarial network(GAN) based method [anand2019fully] can create protein structures by generating pairwise distance maps and decode with a pre-trained downstream neural networks. Another autoencoder base model[eguchi2022ig] focused on immunoglobulins can generate immunoproteins by decoding latent representations. More recently, a model developed can build distance map by iterative refinement and network hallucination[anishchenko2021novo] and repurposed a structure prediction model[yang2020improved] to generate three dimensional protein structures.A recent study that utilized this model demonstrated its ability toe facilitate the design novel enzymes with competitive enzymatic activity[yeh2023novo]. On the other hand, a slew of diffusion based models [watson2023novo, lin2023generating, wu2024protein, ingraham2023illuminating] emerged for structure biology tasks aimed to address challenges presented by previous methods, specifically the ability to flexibly generate diverse and high-quality protein structures. These models have demonstrated success in designing novel monomeric proteins, protein binders, and enzyme active site scaffolds [watson2023novo], with some designs experimentally validated. Despite its success, there are still drawbacks from diffusion based structure models which often demand large amount of compute and hard to incorporate physical constraints into the generative process. It is also worth noting that the success of functional protein design still heavily relies on extensive laboratory screening.
@@
Deep Learning Methods:More recent approaches use deep learning for sequence design, demonstrating superior performance in both computational tests and wet experiments[ingraham2019generative, dauparas2022robust, li2014direct, anand2022protein]. These methods can capture complex, non-linear relationships in protein structure and sequence, and have the ability to learn from large datasets of known protein structures. They are potentially more robust to variations in target structures. However, deep learning methods are often difficult to interpret the basis of their predictions. Additionally, they may struggle with novel protein folds not well-represented in training data[li2014direct, mcpartlon2022deep, gao2022pifold]. Besides, many application specific models are developed for sequence design aimed to enhance performance in domain specific tasks such as antibody design[hoie2024antifold, dreyer2023inverse]. In this study, we will utilize our structure generative model and the iterative design framework in combination with a wide range of inverse folding models to improve the efficiency and robustness of computational protein design pipelines.
Structure-based de novo functional protein design, which involves creating proteins distinct from those exist in the nature based on biophysical principles, has achieved notable successes including the design of novel protein folds[huang2016novo] and the creation of de novo enzymes[rothlisberger2008kemp, yeh2023novo]. Protein redesign and optimization, which involves modifying existing proteins for enhanced stability, altered specificity, or new functions, has also seen substantial advancements [wijma2014computationally]. Although significant effort and progress has been made towards computationally designing novel structure and sequence to achieve new functional proteins, current successes in protein design often require extensive manual input from both biochemistry and computational experts, limiting the accessibility of advanced protein design tools to the broader scientific community and hindering their ability to achieve diverse research goals. This thesis will explore and discuss novel methods for developing a more accessible computational framework, aiming to allow easier access of advanced protein design tools and enable their efficient use for de novo protein design across the life science and computational communities.
@@ 3 Machine learning
KL divergence is non-negative according to the Jensen’s inequality. The log likelihood(evidence) is then lower-bounded by $E_{q_{\phi}}[\log(p_{\theta}(x,z))]-E_{q_{\phi}}[\log(q_{\phi}(z|x))]$, which is referred to as the Evidence Lower Bound(ELBO). We can estimate the gradient of the ELBO with a naive Monte Carlo estimator, however, it is often challenging due to its high variance[kingma2013auto]. Another important observation is that the evidence is a constant and the KL divergence is non-negative, maximizing the ELBO can be seen as simultaneously minimizing the KL divergence between the variational distribution and the true posterior and maximizing the marginalized data likelihood. We can then rewrite the ELBO as
@@ 3 Introduction
In recent years, progress in machine learning and deep learning research has contributed to significant advances for protein modeling such as mutation effect estimation [frazer2021disease, meier2021language, riesselman2018deep], protein function prediction [lai2022accurate, gligorijevic2021structure], and structure prediction [jumper2021highly, baek2021accurate, wang2017accurate].These advancements have also played a part in aspects of the computational protein design problem. For fixed backbone sequence design, a series of deep learning methods have emerged to improve conventional energy based approaches [alford2017rosetta, khatib2011algorithm] by directly incorporating structural information using SE(3)-equivaraint frameworks [mcpartlon2022deep, jing2020learning, hsu2022learning]. An array of recent works have studied the use of generative models for structure generation[lin2021deep, anand2019fully, anand2018generative], however, these methods often generate topological constraints and rely on downstream tools for 3-dimensional structure determination. One of the major difficulties for direct coordinate generative modelling is properly accounting for the rotation and translation equivariance in the target conformation.
@@ 2 Ablation study on input features
In addition, we also compared the short, medium, long range contact accuracy where the predicted contacts are derived from the inter-residue distance of the generated 3-dimensional coordinates as shown in Table 2.2
@@ 3 Conformational decoy sampling for robust protein sequence design
We provide examples in Figure 4(A) where our designed sequence folded more successfully than the native sequence when MSA and template information is not available. Particularly, as shown for PDB:3G67, the native sequence is folded with RMSD = 36.26 while our best decoy designed sequence folded with RMSD = 2.56 with high confidence. We observed that our model can consistently improve designed quality regardless of the size of the target backbone. In Figure 4 (B), we deployed the design pipeline to a set of 211 backbone targets and plotted the TM-score of the predicted structure compared to native sequence(bottom), and sequence designed from the native backbone structure(top). Sequences designed from our pipeline outperform sequences designed from fixed native backbones in 856 out of the 1,016 tested targets(84%).
@@
To further demonstrate our deep structure generative model presents an edge in providing diverse backbone ensembles that can overcome limitation in inverse folding methods relying on single-shot native backbones. Fig. 6(A) illustrates an example where proteinMPNN temperature-based sequence sampling with a native backbone failed to generate structure-conforming sequences, while our structure ensemble approach yielded many successful designs. The advantages of our method is further evidenced in Fig. 6(B), which shows the distribution of sc-TM scores versus pLDDT across all tested targets. Notably, our conformer-based designs (red points) achieved a higher concentration of successful designs in the upper-right quadrant compared to single-shot native backbone templates (blue points). The bar chart below quantifies this advantage, showing that conformer-based designs consistently outperform native backbone designs across various sc-TM cutoffs, with approximately twice the success rate in many cases. Fig. 6(C) further supports this conclusion by showing that our conformer-based designs (blue points) consistently achieve higher pLDDT scores compared to native backbone designs across the majority of tested folds.
@@ 7 Discussion
In the following chapters, I will employ the structural generative model developed and explore de novo protein design and protein optimization through a iterative framework and applying the model to challenging real-world protein design problems and validating designs experimentally. Extending the model to handle multi-chain protein complexes and further antibody design. However, other potential directions for future work includes integrating the structural generative model more tightly with sequence design algorithms, potentially in an end-to-end differentiable framework and exploring the use of the learned structural embeddings for other tasks such as function prediction or protein-protein interaction modeling.
@@ 5 Iterative optimization algorithms
To begin, a randomly sampled sequence pool $Seq_{init}$ of length $L$ is provided. For each iteration, we predetermine the number of templates we select for each design iteration and the number of decoys to generate for each template. Then we set the number of iteration to be $T$. In the design pipeline, Alphafold2 with single sequence mode was used as StructurePrediction and ProteinMPNN was used as the FixedBBDesign algorithm with 1 sequence generated for each decoy. RankPTM is a function that rank a set of prediction structure by their pTM scores.
@@
1:procedure MotifGroundedScaffold($Template_{init},T,N_{decoy},N_{template},MotifSites$)
@@
We initiate our design pipeline with a structure motif $Template_{init}$ of interest as well as the relative infilling among the segments of the motif, for each iteration, we predetermine the number of templates we select for each design iteration and the number of decoys to generate for each template. Then we set the number of iteration to be $T$. In the design pipeline, ESMFold is used as StructurePredictor and ProteinMPNN is used as the FixedBBDesign algorithm with 1 sequence generated for each decoy and the motif sequence were fixed in this step. RankMotifRMSD is a function that rank a set of prediction structure by their the RMSD w.r.t the input motif.
@@ 6 Data
We used the DIPS(Database of Interacting Protein Structures) dataset presented by [townshend2019end] to fine tune our structure generative model for binding surface generation. We capped the size of the complex structure to 500 amino acid to account for memory usage. Then we computed the binding surface with a 10A radius inter-chain contact, that is, any residue that is within 10A to any residue of the other chain. During training, the binding interface of one component of the protein complex is randomly masked.
@@ 2 De novo design of DFHBI activated fluorescent protein
Figure 8: Experimental validation of de novo designed DFHBI-activated fluorescent $\beta$-barrels A) Fluorescence microscope image of de novo designed protein(Left) and the design template mFAP0. B) Fluorescent emission spectras from the lysates of 20 tested designs and the templates with reference buffer. All of the 20 designs were found with detectable fluorescent emission which represents a 100% design success rate. C) Fluorescent emission spectras of purified designed proteins and their respective original templates of 6czi(up) and 6czh(down). Although we did not observe increased fluorescent peak, we saw slight shift in the emission frequency compared to the design template. D)SDS–PAGE of purified designed protein with its respective design templates. All the designed proteins are observed to be smaller than the reference template. E) Protein thermal stability assay results(top), and protein yield comparison results(bottom). We see improvement of thermal stability on 6czh-based designs and comparable thermal performance in 6czi-based designs. For protein yield, all of the 6czh-based designs exhibited improvement over the design reference and 6czi-V has a significant increase over the reference.
We selected the top 20 designs from our computational pipeline for initial experimental screening. These designs were expressed in E. coli(See Methods for details) using standard protocols. Crude cell lysates were then subjected to a fluorescence assay in the presence and absence of DFHBI. This initial screen allowed us to rapidly assess the fluorescence activation of our designs and compare their performance to the design templates. Results from this initial screen revealed that 20 out of 20 designs showed detectable DFHBI-activated fluorescence with a $100\%$ success rate(8.B). All the tested designs have sequence similarity to the design template greater than $74\%$ with the lowest candidate to have $65\%$ sequence identity to the design template as shown in 8.A.
@@
Protein scaffolding is a crucial tool in protein engineering that involves transplanting functional motifs from one protein context to another, often more stable and versatile, protein scaffold[wang2022scaffolding]. This technique has numerous applications in biotechnology and therapeutics, allowing for the creation of proteins with desired functions in optimized structural contexts. Scaffolding can improve protein stability, enhance expression levels, and even modulate the activity of the transplanted motif. To trial our proposed iterative design framework’s ability to embed known functional motif into new protein scaffolds by iteratively co-optimizing structure and sequence around the dedicated motif of interest as seen in 9, we performed an excise on computationally scaffold the PD1/PD-L1 complex and validated our design with in silico binding and folding experiments.
@@
Programmed Cell Death Protein 1 (PD1) is a crucial immune checkpoint receptor that plays a significant role in regulating T-cell responses[sharpe2018diverse]. The interaction between PD1 and its ligands (PD-L1 and PD-L2) is a key target for cancer immunotherapy[sun2018regulation]. However, the use of native PD1 protein or its extracellular domain in therapeutic applications can be challenging due to stability and manufacturing consideration[ganesan2019comprehensive]. In this study, we aimed to use our iterative design framework to scaffold the key binding motif of PD1 onto a more stable protein structure. Our goal was to create a novel protein that maintains PD1’s binding specificity and affinity for its ligands while potentially offering improved biophysical properties.
@@ 4 Structure based Antibody design via conditional CDR inpainting
In Fig11 we observed that the two stage fine tuning significantly improved our model’s ability to accurately in paint the CDRs of interest. Our final fine tuned model $CoordVAE_{ab}$, demonstrates significant improvements across all three CDR regions compared to other models. It consistently achieves the highest TM-scores and pLDDT scores, indicating better overall structural similarity and prediction confidence. Notably, it also maintains the lowest RMSD values across all CDR types, suggesting more accurate reconstruction of the original structures. The performance gap is most significant for CDR3, which is often the most consequential CDR and the hardest to design. These results indicate that our model is exceptionally effective at generating accurate and structurally similar CDR regions, which is crucial for antibody design and engineering.